# What’s the Best Seat in the Game Left, Center, Right?

### v1

@@ Abstract.
Because dice rolls completely determine the game, it lends itself to study using the theory of Markov chains. A Markov chain is a stochastic model in which, at a given state, the probability of reaching another state is fixed and does not depend on previous events. Chutes and Ladders—in which gameplay is determined by the fixed locations of chutes and ladders on the board and spins of the spinner—is another familiar game that can be modeled using Markov chains. For a variety of examples, see [1, 2, 3, 4, 5, 6, 7, 9, 10, 11]. For an analysis of a variant of Left, Center, Right that does not use Markov chains, see [13].
@@ 1. An analysis using Markov chains
To study Left, Center, Right mathematically, we first define the state of the game. If there are $n$ players, we can represent the game state by an $(n+1)$-tuple of non-negative integer values, $(c_{1},c_{2},\ldots,c_{n};p)$, where $c_{i}$ is the number of chips in front of player $i$ and $p$ indicates whose turn it is.
@@
Once we know the transition probabilities between all the states, we obtain a directed graph with probabilities on each edge. Playing the game corresponds to beginning at the state $(3,3,\ldots,3;1)$ and traversing the graph, making decisions based on the probabilities on the edges. The game ends when we get to a state $(c_{1},c_{2},\ldots,c_{n};p)$ in which only one of the $c_{i}$ is nonzero. In the directed graph, such vertices have a loop from the vertex to itself labeled with a probability of 1.
@@
Give none $\frac{1}{2}$ $\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4}$ $\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{8}$
@@
Note also that if it is player $i$’s turn but they have no chips, then with probability 1, the game transitions to the next state (chip counts remain the same, and it is player $(i+1)$’s turn to roll).
@@
where $Q$ is a $t\times t$ matrix, $R$ is an $t\times a$ matrix, $\mathbf{0}$ is an $a\times t$ matrix of 0’s, and $I_{a}$ is the $a\times a$ identity matrix.
@@
$\mathbf{w}=N\left(\begin{smallmatrix}1\\ 1\\ \vdots\\ 1\end{smallmatrix}\right)=\left(\begin{smallmatrix}w_{1}\\ w_{2}\\ \vdots\\ w_{r}\end{smallmatrix}\right)\text{ and }\mathbf{w}_{\text{sq}}=\left(\begin{smallmatrix}w_{1}^{2}\\ w_{2}^{2}\\ \vdots\\ w_{r}^{2}\end{smallmatrix}\right).$
@@
$\left(2N-I_{r}\right)\mathbf{w}-\mathbf{w}_{\text{sq}}.$
@@
For instance, we computed that the average game length for three players is 18.9 rolls with a standard deviation of 8.1.
@@
is the probability of the game ending in absorbing state $j$ when starting from transient state $i$. Assuming, again, that we begin at vertex 1, then the sum of all the entries $(1,j)$ in which $j$ corresponds to a victory for player $p$ is the probability that player $p$ wins.
@@ 2. Monte Carlo simulation
Suppose we play $R$ games in our Monte Carlo simulation. When we do, we obtain an approximation $\overline{x}$ for the average number of turns to win or the fraction of the games won by one of the players. By the law of large numbers, the larger $R,$ the closer $\overline{x}$ is to the desired quantity. Moreover, by the central limit theorem, the results of such Monte Carlo simulations follow a roughly normal distribution centered around the desired value. In this scenario, the Monte Carlo error, $\sigma_{R}$, is the standard deviation of the Monte Carlo estimator taken across hypothetical repetitions of the simulation, where each simulation consists of $R$ replications. (See [8] or [12, §3.4], for instance.) If $\hat{\sigma}$ is the calculated standard deviation obtained from the $R$ games,
@@
Obtaining multiple decimal places of accuracy for the expected game length is unnecessary, but we do require a high degree of accuracy when trying to see which players are most and least likely to win.
Suppose we’ve played $R$ simulated games, and player $k$ wins $w$ of them. Then, $\overline{x}=w/R$. We can use the same technique to compute the Monte Carlo error as we did for the average game length, but in this case, $x_{i}$ is either 1 or 0, depending on whether the player wins or does not win game $i$, respectively. We see that
$\sigma_{R}$ $\approx\sqrt{\frac{1}{R^{2}}\sum_{i=1}^{R}\left(x_{i}-\overline{x}\right)^{2}}$
$=\sqrt{\frac{1}{R^{2}}\left(w\left(1-\overline{x}\right)^{2}+(R-w)\left(0-\overline{x}\right)^{2}\right)}$
$=\sqrt{\frac{1}{R^{2}}\left(w\left(1-\frac{w}{R}\right)^{2}+(R-w)\left(\frac{w}{R}\right)^{2}\right)}$
$=\sqrt{\frac{Rw-w^{2}}{R^{3}}}$
For instance, in our Monte Carlo simulation of five-player games, player 1 won $w=19,375,770$ of the $R=100,000,000$ games. So $\overline{x}=0.1937577.$ Then,
$2\sigma_{R}\approx 2\sqrt{\frac{(10^{8})(19375770)-(19375770)^{2}}{(10^{8})^{3}}}\approx 0.000079048$
@@ 4. Discussions and conclusions
The game begins in an unusual fashion. The first player will have three chips on their first turn, but the rest will, on average, have more than three on their first turn since the neighbor on their right may have passed them some. Moreover, the last player, player $n$, may receive chips from two players before they roll—from players 1 and $n-1$. For more consistency, one could make one simple rule change: When player 1 rolls for the first time, they treat any R’s as H’s. That is, they hold the chips they would otherwise pass to player $n$. We will call this the modified start rule.
@@
Still, if each of the eight players starts with \$30, player 7’s expected winnings of \$$\$30\cdot 8\cdot 0.13-\$30=\$1.12$ is quite a bit better than player 2’s expected loss of \$$\$30-\$30\cdot 8\cdot 0.12=\$1.20$.

### v2

@@ Abstract.
Because dice rolls completely determine the game, it lends itself to study using the theory of Markov chains. A Markov chain is a stochastic model in which, at a given state, the probability of reaching another state is fixed and does not depend on previous events. Chutes and Ladders—in which gameplay is determined by the fixed locations of chutes and ladders on the board and spins of the spinner—is another familiar game that can be modeled using Markov chains. For a variety of examples, see [1, 2, 3, 4, 5, 6, 7, 9, 10, 11]. For an analysis of a variant of Left, Center, Right that does not use Markov chains, see [14]. No previous exposure to Markov chains is required to follow the arguments in this article, but to learn more about this area of probability, see, for instance, [12].
@@ 1. An analysis using Markov chains
To study Left, Center, Right mathematically, we first define the state of the game. If there are $n$ players, we can represent the game state by an $(n+1)$-tuple of non-negative integer values, $(c_{1},c_{2},\ldots,c_{n};j)$, where $c_{i}$ is the number of chips in front of player $i$, and $j$ indicates whose turn it is.
@@
Once we know the transition probabilities between all the states, we obtain a directed graph with probabilities on each edge. Playing the game corresponds to beginning at the state $(3,3,\ldots,3;1)$ and traversing the graph, making decisions based on the probabilities on the edges. The game ends when we get to a state $(c_{1},c_{2},\ldots,c_{n};j)$ in which only one of the $c_{i}$ is nonzero. In the directed graph, such vertices have a loop from the vertex to itself labeled with a probability of 1.
@@
No L, C, or R rolls $\frac{1}{2}$ $\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4}$ $\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{8}$
@@
Note also that if it is player $i$’s turn but they have no chips, then with probability 1, the game transitions to the next state; chip counts remain the same, and it is player $(i+1)$’s turn to roll.
@@
where $Q$ is a $t\times t$ matrix, $R$ is a $t\times a$ matrix, $\mathbf{0}$ is an $a\times t$ matrix of 0’s, and $I_{a}$ is the $a\times a$ identity matrix.
@@
$\mathbf{w}=N\left(\begin{smallmatrix}1\\ 1\\ \vdots\\ 1\end{smallmatrix}\right)=\left(\begin{smallmatrix}w_{1}\\ w_{2}\\ \vdots\\ w_{t}\end{smallmatrix}\right)\text{ and }\mathbf{w}_{\text{sq}}=\left(\begin{smallmatrix}w_{1}^{2}\\ w_{2}^{2}\\ \vdots\\ w_{t}^{2}\end{smallmatrix}\right).$
@@
$\left(2N-I_{t}\right)\mathbf{w}-\mathbf{w}_{\text{sq}}.$
@@
For instance, we computed that the average game length for three players is 18.9 rolls with a standard deviation of 8.1 rolls.
@@
is the probability of the game ending in absorbing state $j$ when starting from transient state $i$. Assuming, again, that we begin at vertex 1, then the sum of all the entries $(1,j)$ of $B$ in which $j$ corresponds to a victory for player $p$ is the probability that player $p$ wins.
@@ 2. Monte Carlo simulation
Suppose we play $R$ games in our Monte Carlo simulation. When we do, we obtain an approximation $\overline{x}$ for the average number of turns to win or the fraction of the games won by one of the players. By the law of large numbers, the larger $R,$ the closer $\overline{x}$ is to the desired quantity. Moreover, by the central limit theorem, the results of such Monte Carlo simulations follow a roughly normal distribution centered around the desired value. In this scenario, the Monte Carlo error, $\sigma_{R}$, is the standard deviation of the Monte Carlo estimator taken across hypothetical repetitions of the simulation, where each simulation consists of $R$ replications. (See [8] or [13, §3.4], for instance.) If $\hat{\sigma}$ is the calculated standard deviation obtained from the $R$ games,
@@
Obtaining multiple decimal places of accuracy for the expected game length is unnecessary, but we do require a high degree of accuracy when trying to see which players are most and least likely to win. While we could use the Monte Carlo error formula to compute the error, there is a more straightforward method.
Suppose we’ve played $R$ simulated games, and player $k$ wins $w$ of them. Then $\hat{p}=w/R$ is the sample proportion of games won. If $p$ is the true proportion, the standard error is
$\sigma_{\hat{p}}$ $=\sqrt{\frac{p(1-p)}{R}}$
$\approx\sqrt{\frac{\hat{p}(1-\hat{p})}{R}}$
$=\sqrt{\frac{1}{R}\left(\frac{w}{R}\right)\left(1-\frac{w}{R}\right)}$
$=\sqrt{\frac{Rw-w^{2}}{R^{3}}}=\hat{\sigma}_{\hat{p}}.$
Moreover, the Central Limit Theorem tells us that $(\hat{p}-p)/\hat{\sigma}_{\hat{p}}$ converges in distribution to a standard normal random variable.
For instance, in our Monte Carlo simulation of five-player games, player 1 won $w=19,375,770$ of the $R=100,000,000$ games. So $\hat{p}=0.1937577.$ Then,
$2\hat{\sigma}_{\hat{p}}=2\sqrt{\frac{(10^{8})(19375770)-(19375770)^{2}}{(10^{8})^{3}}}\approx 0.000079048$
@@ 4. Discussions and conclusions
The game begins in an unusual fashion. The first player will have three chips on their first turn, but the rest will, on average, have more than three on their first turn since the neighbor on their right may have passed them some. Moreover, the last player, player $n$, may receive chips from two players before they roll—from players 1 and $n-1$. For more consistency, one could make one simple rule change: When player 1 rolls for the first time, they treat any L’s as H’s. That is, they hold the chips they would otherwise pass to player $n$. We will call this the modified start rule.
@@
Still, if each of the eight players starts with \$30, player 7’s expected winnings of \$$\$30\cdot 8\cdot 0.13-\$30=\$1.20$ is quite a bit better than player 2’s expected loss of the same amount.